OpenStudy (anonymous):

its everybody again! and everybody needs help (again) (/ is a fraction bar) 18 divided by 4/5 =??? plz help best explanation will earn a medal!

OpenStudy (anonymous):

@Romana101

OpenStudy (anonymous):

anyone??

OpenStudy (anonymous):

yey

OpenStudy (igreen):

Flip the fraction and multiply. 18 * 5/4

OpenStudy (igreen):

\(\dfrac{18 \times 5}{4}\) Simplify

OpenStudy (anonymous):

18 x 5 = 90 / 4 = 22.5

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank you!

OpenStudy (igreen):

Yep, that's our answer.